# Observation of Kolmogorov turbulence due to multiscale vortices in dusty plasma experiments

### v1

@@ Abstract
We report the experimental observation of fully developed Kolmogorov turbulence originating from self-excited vortex flows in a three-dimensional (3D) dust cloud. The characteristic -5/3 scaling of three-dimensional Kolmogorov turbulence is universally observed in both the spatial and temporal power spectra. Additionally, the 2/3 scaling in the second-order structure function further confirms the presence of Kolmogorov turbulence. We also identified a slight deviation in the tails of the probability distribution functions for velocity gradients. The dust cloud formed in the diffused region away from the electrode and above the glass device surface in the glow discharge experiments. The dust rotation was observed in multiple experimental campaigns under different discharge conditions at different spatial locations and background plasma environments.
@@ I Introduction
Laboratory dusty (complex) plasmas comprise of electrons, ions, neutral gas particles, and highly charged dust particles (Melzer et al., 2019; Shukla and Mamun, 2015). Dust particles generally have sizes up to tens of $\mu$m and typically attains a high negative charge of $10^{4}$e due to the high mobility of electrons. However, under atypical conditions, the dust particles can also acquire positive charges transiently or stably, as reported in various afterglow dusty plasma experiments (Chaubey et al., 2021; Samarian et al., 2001). Their dynamics are slow (a few tens of Hz) compared to electron and ion timescales. The length scales usually ranges from a few millimeters to a few centimeters (Melzer et al., 2019). The slow dynamics and convenient length scales render it an ideal experimental platform for accurately monitoring complex physics phenomena. Some examples are phase transition, longitudinal and transverse waves, solitons, shocks, vortices, and complex flows (Bailung et al., 2020; Bandyopadhyay et al., 2008; Jaiswal et al., 2016; Choudhary et al., ; Pramanik et al., 2002; Heinrich et al., 2009; Arora et al., 2021; Kaur et al., 2015a; Bose et al., 2019).
Dusty plasma experiments with precision imaging diagnostics enable the study of complex flows. It includes two and three-dimensional laminar, sheared, rotating, chaotic, and turbulent flows (Nosenko and Goree, 2004; Pramanik et al., 2002; Jaiswal et al., 2016; Gavrikov et al., 2005; Choudhary et al., 2020). These low Reynolds number flows can be classified as either hydrodynamic or visco-elastic. The classification depends on the interaction strength of the strongly charged dust particles relative to their thermal energy in the plasma. Dusty plasma flows are often observed in experiments due to the interplay of various forces acting on dust particles. These flows can be tracked from particle scales to hydrodynamic scales Feng et al. (2007); Thomas et al. (2010); Boessé et al. (2004). In this study, we observed the turbulent flow caused by the dust vortex, particularly within the diffuse region of a DC glow discharge plasma. We found the exact Kolmogorov power law scaling in both spatial and temporal domains of the energy spectrum, thus demonstrating Kolmogorov turbulence in afterglow dusty plasma conditions for the first time.
Dusty plasma experiments provide a unique platform for studying turbulent flows across both macroscopic (system size) and microscopic (average interparticle separation) scales. This precision is achieved by tracking individual dust particles’ trajectories, even in the liquid phase. These experiments serve as an alternative to Navier-Stokes equation-based numerical models, which depend on the continuum approximation. Such models are computationally expensive because they require numerous grid points to resolve scales close to the particle level (Marion and Temam, 1998; Wani et al., 2023).
@@
For dusty plasmas, Schwabe et al.Schwabe et al. (2018a) experimentally demonstrated a dual cascade in the spectrum. They found that $E(k)\propto k^{-5/3}$ for $k<k_{l}$ and $E(k)\propto k^{-3}$ for $k>k_{l}$, where $k_{l}$ is an intermediate forces wavenumber. Their results are consistent with theoretical predictions of two-dimensional forced turbulence (Kraichnan, 1973, 1967, 1974). Tsai and Lin I (Tsai and Lin-I, 2014) demonstrated that the dust acoustic wave turbulence exhibits a frequency spectrum proportional to $f^{-2.6}$. Using kinetic numerical simulations, Joshi et al. (Joshi et al., 2024) demonstrated turbulent mixing due to the vortex motion in the dust flow past the obstacle. Their results follow the hydrodynamic Kolmogorov turbulence scaling of $-5/3$ in their energy spectrum in both the time and spatial domains. While all previous works report wave turbulence, this work is the first to demonstrate fully developed Kolmogorov turbulence in dusty plasma due to multi-scale vortex flow. The vortex flow appears stationary at the largest scale of energy infusion. We could show the scaling of the energy cascade $-5/3$ in the wave-vector and frequency domains.
In this paper, we report a few large-scale vortices and small-scale eddies in a dust cloud. It levitates in the diffused plasma far from the electrode in a DC discharge. The dust cloud remains stable for several hours due to the balance of electrostatic $(q\vec{E})$ and gravitational $(m\vec{g})$ forces. This stability is maintained as long as an external source continuously provides energy and the dust cloud resides in an energy bath of background plasma. The dust cloud is shaped in three dimensions, but the image analysis and particle image velocimetry (PIV) are carried out based on the two-dimensional imaging of the data. We estimate a strong two-order-of-magnitude background plasma density variation between the electrode and the diffused plasma region. Reported vortex observations are highly reproducible in SPD-I device (Sharma et al., 2023). We presented the turbulence features in the rotating dust cloud of the diffused plasma region. Understanding the flow dynamics in diffused plasma regions is essential due to their relevance to the interstellar universe and the lunar surface, which have minor density variations. We observed a Kolmogorov $k^{-5/3}$ energy scale in the frequency and spatial domains of the power spectrum. To further support our results, we have demonstrated the second-order structure functions following Kolmogorov law and the probability distribution function (PDF) of velocity gradients.
@@ II Experimental details
Figure 1: The schematic of SPD-I with a laser sheet in the X-Z plane and a camera lens positioned perpendicularly to capture images. These images are later used for PIV analysis to determine velocity fields in the X and Z directions. Detailed device dimensions can be found in Sharma et al. (2023).
@@
The experiments are performed on the newly commissioned laboratory plasma device Shivalik Plasma Device-I (SPD-I) Sharma et al. (2023). This plasma chamber is a horizontally laid (Z-axis) glass cylinder with two asymmetric electrodes kept vertically in a parallel plate arrangement, as shown in Fig. 1. The glass chamber is electrically floating using Teflon connectors, with the top electrode as an anode and the bottom grounded electrode as a cathode. In the present experiments, we separated both electrodes by 7.5 cm. Figure 1 depicts this DC discharge-based experimental set-up. The details on configuration, including the device dimensions, electrodes, and diagnostics, are in Sharma et al. (2023). The working gas Argon was inserted in the vacuum chamber from a port next to the pumping port to avoid any directed flow of neutral gas in the experimental region of interest, as doing so allows only the diffused gas to enter the system (Kaur et al., 2016). The vacuum vessel was purged with Argon gas multiple times before producing the discharge. At each purging cycle Argon was inserted at high pressure ( $\sim 5$ mbar) for at least 10 minutes and then pumped down to low pressures. The plasma was formed in an Argon gas background with a working pressure of 0.12 mbar.
@@
Polydispersive kaolin particles, ranging in size from 4 to 8 $\mu$m, dispersed over the cathode before creating the vacuum. A slow initial pumping was performed to prevent the Kaolin particles from being forced out of the device. These dust particles over the cathode and some of its floating fraction get charged as soon as plasma creation occurs. A void quickly forms over the cathode because of the competition between the ion drag force and the radial electric field. Almost all floating dust particles are expelled from the electrode during the void formation. However, these dust particles then form a three-dimensional (3D) cloud over the extended sheath on the glass surface. The dust cloud is formed in the diffused plasma region far from the grounded cathode. The 3D cloud forms possibly for two reasons: First, because of polydispersive Kaolin particles, and second, due to the 3D potential well formation in the diffused plasma region. Three forces may contribute to this potential well formation. 1) A radially outward force on dust due to void formation over the cathode. 2) The force due to the lower glass surface’s sheath electric field. 3) The force due to the electric field profiles of the end-flange sheath. The second reason is the predominant cause of the 3D dust cloud in the present experiment. This is because the 3D nature remains with monodispersive MF particles in the SPD-I device Sharma et al. (2023). In the present work, all of the campaigns have been performed using Kaolin microparticles.
@@
A green LASER of wavelength 532 nm with a variable power capacity of up to 300 mW illuminates the dust particles. A cylindrical lens is placed before the LASER to convert it into a thin sheet of thickness 150 $\mu$m. A high-speed camera captures the illuminated dust particles at a 90-degree angle to the LASER sheet (X-Z plane). We have used two different cameras. The PCO Edge 5.5 is utilized for recording up to 100 fps with a resolution of 2560 $\times$ 2180 pixels, chosen for its superior spatial resolution. The Phantom v2512 records up to 1000 fps with a resolution of 1064 $\times$ 1064 pixels, selected to achieve a superior temporal data profile. Particle Image Velocimetry (PIV) has served as the primary tool for extracting velocity fields from the recorded image sequences (detailed in subsection  II.2). Particle positions and velocities are also extracted through particle tracking over time for at least two experimental datasets. We used these particle data to validate PIV results and predict the static structural properties of dust clouds.
@@ II.1 Datasets and plasma parameters
We conducted multiple campaigns to study flows and turbulence to ensure the reproducibility and consistency of the observed dust rotation phenomena. A total of six independent experimental campaigns were conducted at the same working pressure of 0.12 mbar, as listed in Table 1. The schematic in Fig. 2 shows tentative locations of dust rotation from different campaigns. The hemispherical shaded region represents the dust void. In each campaign, we observed dust rotation with variations in flow characteristics. Specifically, I, II, III and VI form in reasonably different plasma environments. We anticipate an interplay of different set of forces Kaur et al. (2015b); Chai and Bellan (2016); Vaulina et al. (2003a); Choudhary et al. (2018); Samarian et al. (2002) to be responsible for individual vortices in the extended diffused region. The suggestive reason is that the left edge of the cloud faces the cathode, and the right edge is far from the plasma source and is towards the glass flange boundary. It is interesting to note that in all these conditions, the nature of turbulence follows the standard forced Kolmogorov turbulence predictions.
Table 1: Campaign details: The pressure is kept at $P=0.12$ mbar for all cases. The table provides the discharge voltage, $V_{d}$ (in volts), and the tentative cloud span, $S_{xz}$. It also includes their horizontal distance from the cathode, $d_{hc}$, and the vertical distance from the glass surface, $d_{vg}$.
@@
III 536 $12\times 35$ 50 1
IV 572 $7\times 55$ 55 1.1
V 538 $11\times 35$ 50 1
@@
Epstein drag coefficient (Epstein, 1924; Pustylnik et al., 2006), $\gamma_{ep}$ $8~{}\mathrm{s^{-1}}$
Figure 3: (a) The temperature of dust-charged particles is determined by analyzing the variance of the fitted Maxwellian distribution to particle distribution functions. (b) Particle positions from the 2D image (XZ plane) observed in Campaign No. VI. The particle data was extracted using ImageJ software Schindelin et al. (2012). The region for data extraction was taken from the no-flow equilibrium region, reflecting only thermal motions.
@@
Figure 4: The radial distribution functions, $g(r)$ obtained from the particle data (blue solid and magenta dashed lines) from two experimental datasets.
@@
Table 2 enlists parameters for the background plasmas and the dust cloud. The electron temperature, $T_{e}$, and the plasma density, $n_{e,i}$ (in quasineutral plasma), were measured using the single Langmuir probe above the electrode and in the diffused plasma region in the absence of dust Bose et al. (2017). Over the cathode, $T_{e}$ ranges $1.5-4$ eV from the center to the edge. In the diffused region where dusty plasma typically forms, $T_{e}$ is relatively high, ranging between 3 and 7 eV at different locations within the domain span. The plasma density varies by up to two orders of magnitude from the electrode to the diffused region of interest, as shown in the table.
@@
We used a microscope to obtain the kaolin particles’ size distribution. Most particles ranged from 5 to 8 $\mu$m, although there were a few outliers with sizes as large as 50 $\mu$m. Based on the known mass density of kaolin used in the experiments, the typical mass of dust particles ranged from $1.6\times 10^{-13}$ to $6.7\times 10^{-13}$ kg. Out of the six campaigns, we could track particle positions with reasonable accuracy using the ImageJ software Schindelin et al. (2012) for three campaigns. We used the particle data from Fig. 3(b) for 100 successive frames and a particle tracking algorithm to gather particle velocities (in mm/s). Determining the accurate average interparticle separation, $a$, for dust clouds composed of particles of varying sizes and experiencing different spatial plasma environments is challenging. However, we calculated $a$ from regions near vortex motion with no flow but only thermal motion. We obtained an estimated value of $a\approx 224~{}\mu$m and a typical dust density, $n_{d}\approx 2.3\times 10^{4}$ cm-3 in the diffused plasma region. Figure 3(b) depicts one such particle arrangement (from Campaign VI) we used to estimate $a$.
@@
The typical charge on dust particles ranges from $2$ to $5\times 10^{4}$$e$. The charge is calculated using the collisionless orbital motion limited (OML) model, based on the floating potential on a spherical dust particle. We plotted the velocity histogram in Fig. 3(a) and fitted it with the theoretical Maxwellian speed distribution. A comparison between experimental data and theoretical fitting gives dust temperature ranges from 0.5 to 1.8 eV. We note that the dust temperature is calculated using a two-dimensional velocity field, so it will differ in a three-dimensional scenario. We also calculated the 2D radial distribution function (RDF) for two datasets (II and IV). The region chosen for this calculation does not have bulk flow, and the dynamics are thermal. The RDF in Fig. 4 with a small first peak shows the equilibrium dust dynamics in a fairly liquid regime.

### v2

@@ Abstract
We report the experimental observation of fully developed Kolmogorov turbulence originating from self-excited vortex flows in a three-dimensional (3D) dust cloud. The characteristic -5/3 scaling of 3D Kolmogorov turbulence is consistent in both the spatial and temporal energy spectra within a statistical variation of experimental data. Additionally, the 2/3 scaling in the second-order structure function further supports the presence of Kolmogorov turbulence. We also identified a slight deviation in the tails of the probability distribution functions for velocity gradients, a reflection of intermittency. The experiment showed the formation of a dust cloud in the diffused plasma region away from the electrodes. The dust rotation was observed in multiple experimental campaigns under different discharge conditions at different spatial locations and background plasma environments.
@@ I Introduction
Laboratory dusty (complex) plasmas comprise of electrons, ions, neutral gas particles, and highly charged dust particles (Melzer et al., 2019; Shukla and Mamun, 2015). Dust particles generally have sizes up to tens of $\mu$m and typically attain a high negative charge of $10^{4}$e due to the high mobility of electrons. However, under atypical conditions, the dust particles can also acquire positive charges transiently or stably, as reported in various afterglow dusty plasma experiments (Chaubey et al., 2021; Samarian et al., 2001). Their dynamics are slow (a few tens of Hz) compared to electron and ion timescales. The length scales usually ranges from a few millimeters to a few centimeters (Melzer et al., 2019). The slow dynamics and convenient length scales render it an ideal experimental platform for accurately monitoring complex physics phenomena. Some examples are phase transition, longitudinal and transverse waves, solitons, shocks, vortices, and complex flows (Bailung et al., 2020; Bandyopadhyay et al., 2008; Jaiswal et al., 2016; Choudhary et al., ; Pramanik et al., 2002; Heinrich et al., 2009; Arora et al., 2021; Kaur et al., 2015a; Bose et al., 2019).
Dusty plasma experiments with precision imaging diagnostics enable the study of complex flows. It includes two and three-dimensional laminar, sheared, rotating, chaotic, and turbulent flows (Nosenko and Goree, 2004; Pramanik et al., 2002; Jaiswal et al., 2016; Gavrikov et al., 2005; Choudhary et al., 2020). These low Reynolds number flows can be classified as either hydrodynamic or visco-elastic. This classification depends on the interaction strength of the charged dust particles relative to their thermal energy in the plasma. Dusty plasma flows are often observed in experiments due to the interplay of various forces acting on dust particles. These flows can be tracked from particle scales to hydrodynamic scales Feng et al. (2007); Thomas et al. (2010); Boessé et al. (2004). In this study, we observed the turbulent flow caused by the dust vortex, particularly within the diffuse region of a DC glow discharge plasma. We found the alignment with Kolmogorov power law scaling in both spatial and temporal domains of the energy spectrum, thus demonstrating Kolmogorov turbulence in afterglow-like dusty plasma conditions.
Dusty plasma experiments provide a unique platform for studying turbulent flows across both macroscopic (system size) and microscopic (average interparticle separation) scales. This precision is achieved by tracking individual dust particles’ trajectories, even in the liquid phase. These experiments serve as an alternative to Navier-Stokes equation-based numerical models, which depend on the continuum approximation. Such models are computationally expensive because they require numerous grid points to resolve scales close to the particle level (Marion and Temam, 1998; Wani et al., 2024).
@@
For dusty plasmas, Schwabe et al.Schwabe et al. (2018a) experimentally demonstrated a dual cascade in the spectrum. They found that $E(k)\propto k^{-5/3}$ for $k<k_{l}$ and $E(k)\propto k^{-3}$ for $k>k_{l}$, where $k_{l}$ is an intermediate forces wavenumber. Their results are consistent with theoretical predictions of two-dimensional forced turbulence (Kraichnan, 1973, 1967, 1974). Tsai and Lin I (Tsai and Lin-I, 2014) demonstrated that the dust acoustic wave turbulence exhibits $f^{-2.6}$ scaling in a frequency spectrum. Using kinetic numerical simulations, Joshi et al. (Joshi et al., 2024) demonstrated turbulent mixing due to the vortex motion in the dust flow past the obstacle. Their results follow the hydrodynamic Kolmogorov turbulence scaling of $-5/3$ in their energy spectrum in both the time and spatial domains. Using molecular dynamics simulation, Wani et al. (2024) demonstrated the turbulence in a dusty plasma system using Rayleigh-Taylor instability as an initial condition. Their spectra align with the Bolgiano-Obukho theory, a typical two-dimensional Rayleigh-Taylor turbulence. While all previous experimental works report wave turbulence, this work is the first to demonstrate fully developed Kolmogorov turbulence in dusty plasma due to multi-scale vortex flow. The vortex flow appears stationary at the largest scale of energy infusion. We could show the scaling of the energy cascade close to $-5/3$ in the wavevector and frequency domains.
In this paper, we report a few large-scale vortices and small-scale eddies in a dust cloud. It levitates in the diffused plasma far from the electrode in a DC discharge. The dust cloud remains stable for several hours due to the balance of electrostatic $(q\vec{E})$ and gravitational $(m\vec{g})$ forces. This stability is maintained as long as an external source continuously provides energy, and the dust cloud resides in an energy bath of background plasma. The dust cloud is 3D in shape. However, the image analysis and Particle Image Velocimetry (PIV) are carried out on the basis of the two-dimensional imaging of the cloud. We estimate two-order-of-magnitude background plasma density variations between the electrode and the diffused plasma region. The observed vortex is reproducible on the SPD-I device (Sharma et al., 2023). We presented the turbulence features in the rotating dust cloud of the diffused plasma region. Understanding the flow dynamics in diffused plasma regions is interesting due to their relevance to the interstellar universe and the lunar surface, which have minor density variations. We observe a scaling consistent with Kolmogorov’s ${-5/3}$ in the frequency and spatial domains of the energy spectra. Also, in all energy spectra calculations, the kinetic energy is computed using only the fluid velocity components without incorporating the masses of fluid elements. To further support our results, we have demonstrated the second-order structure functions following the Kolmogorov law and the Probability distribution function (PDF) of velocity gradients.
@@ II Experimental details
Figure 1: The schematic of SPD-I shows a laser sheet in the X-Z plane and a camera lens positioned perpendicularly to capture images. These images are used for PIV analysis to determine velocity fields in the X and Z directions. For detailed device dimensions, see Sharma et al. (2023).
@@
The experiments are performed on a newly commissioned laboratory plasma device, Shivalik Plasma Device-I (SPD-I) Sharma et al. (2023). This plasma chamber is a horizontally laid (Z-axis) glass cylinder with two asymmetric electrodes kept vertically in a parallel plate arrangement, as shown in Fig. 1. The glass chamber is electrically floating using Teflon connectors, with the top electrode as an anode and the bottom grounded electrode as a cathode. In the present experiments, we separated both electrodes by 7.5 cm. Figure 1 depicts this DC discharge-based experimental set-up. The details on configuration, including the device dimensions, electrodes, and diagnostics, are in Sharma et al. (2023). We introduced Argon, the working gas, into the vacuum chamber through a port adjacent to the pumping port to prevent any directed flow of neutral gas into the experimental region of interest. This approach ensures that only diffused gas enters the system (Kaur et al., 2016). We repeatedly purged the device with Argon gas at high pressures ($\sim 5$ mbar) for at least 10 minutes before maintaining working pressure to conduct the experiment. The plasma was formed in an Argon gas background at a low working pressure of 0.12 mbar.
@@
Polydispersive kaolin particles, ranging in size from 4 to 8 $\mu$m, are dispersed over the cathode before creating the vacuum. We performed a slow initial pumping to ensure the Kaolin particles remained inside the device. As soon as plasma creation takes place, the dust particles covering the cathode, along with some of its floating fraction, become charged. A void quickly forms over the cathode because of the competition between the ion drag force and the radial electric field. During the void formation, the electrode expels almost all floating dust particles. However, these dust particles then form a 3D cloud in the diffused plasma region over the extended sheath on the glass surface. The 3D cloud forms, possibly for two reasons: first, because of polydispersive Kaolin particles, and second, because of the 3D potential well formation in the diffused plasma region. The second reason is the predominant cause of the 3D dust cloud in the present experiment. We validated it using monodispersive MF particles in the SPD-I device Sharma et al. (2023) and a similar 3D dust cloud formed. Three forces may contribute to this potential well formation. 1) A radially outward force on dust due to void formation over the cathode. 2) The force generated by the lower glass surface’s sheath electric field. 3) The force emanating from the electric field profiles of the end-flange sheath.
@@
A green LASER with a 532 nm wavelength and a variable power capacity of up to 300 mW illuminates the dust particles. A cylindrical lens is placed before the LASER to convert it into a thin sheet of thickness 150 $\mu$m. A high-speed camera captures the illuminated dust particles at a 90-degree angle to the LASER sheet (X-Z plane). We have also made a few observations by keeping the LASER sheet in the Y-Z plane to support the three-dimensional nature of the dynamics. We have used two different cameras. The PCO Edge 5.5 is utilized for recording up to 100 fps with a resolution of 2560 $\times$ 2180 pixels, chosen for its superior spatial resolution. The Phantom v2512 records up to 1000 fps with a resolution of 1064 $\times$ 1064 pixels, selected to achieve a superior temporal data profile. PIV has served as the primary tool for extracting velocity fields from the recorded image sequences (detailed in subsection  II.2). We also extract particle positions and velocities by tracking particles over time for at least two experimental datasets. We used these particle data sets to validate PIV results and predict the static structural properties of dust clouds.
@@ II.1 Datasets and plasma parameters
We conducted multiple campaigns to study flows and turbulence to ensure the reproducibility and consistency of observed dust rotation phenomena. A total of six independent experimental campaigns were conducted at the same working pressure of 0.12 mbar, as listed in Table 1. All of these campaigns have been performed using polydispersive Kaolin microparticles. The schematic in Fig. 2 shows tentative dust rotation locations from different campaigns. These campaigns successfully captured the X-Z plane of the dust cloud, ensuring that the laser sheet remained at the center of the glass flange at location $y=0$. The hemispherical, shaded region represents the dust void. In each campaign, we observed dust rotation with variations in flow characteristics. Specifically, I, II, III, and VI form in reasonably different plasma environments. We anticipate an interplay of different sets of forces Kaur et al. (2015b); Chai and Bellan (2016); Vaulina et al. (2003a); Choudhary et al. (2018); Samarian et al. (2002) to be responsible for individual vortices in the extended diffused region. It is interesting to note that in all these conditions, the nature of turbulence almost follows the standard forced Kolmogorov theory.
Table 1: Campaign details: The table provides the discharge voltage, $V_{d}$, and the tentative cloud span, $S_{xz}$. It also includes their horizontal distance from the cathode, $d_{hc}$, and the vertical distance from the glass surface, $d_{vg}$.
@@
III 565 $7\times 55$ 55 1.1
IV 536 $12\times 35$ 50 1
V 538 $11\times 35$ 60 1
@@
Epstein damping rate (Epstein, 1924; Pustylnik et al., 2006), $\gamma_{ep}$ $8~{}\mathrm{s^{-1}}$
Figure 3: (a) Normalized probability distribution function of particle velocities fitted with the Gaussian distribution. The charged dust particle temperature is obtained from full-width half maxima of fitted Gaussian. (b) Particle positions in the YZ plane from campaign V. The extraction region corresponds to a visually no-flow equilibrium zone, capturing dominantly the thermal motion.
@@
Figure 4: The radial distribution functions, $g(r)$ obtained from the particle data for two datasets.
@@
Table 2 enlists parameters for the background plasmas and the dust cloud. The electron temperature, $T_{e}$, and the plasma density, $n_{e,i}$ (in quasineutral plasma), were measured using the single Langmuir probe above the electrode and in the diffused plasma region in the absence of dust Bose et al. (2017). Over the cathode, $T_{e}$ ranges $1.5-4$ eV from the center to the edge. In the diffused region where dusty plasma typically forms, $T_{e}$ is relatively high, ranging from 3 - 7 eV at different locations within the domain span. The plasma density varies by up to two orders of magnitude from the electrode to the diffused region of interest, as shown in the table.
@@
We used a microscope to obtain the kaolin particles’ size distribution. Most particles ranged from 5 to 8 $\mu$m, although there were a few outliers with sizes as large as 50 $\mu$m. Based on the known mass density of kaolin used in the experiments, the typical mass of dust particles ranged from $1.6\times 10^{-13}$ to $6.7\times 10^{-13}$ kg. Out of the six campaigns, we could track particle positions with reasonable accuracy using the ImageJ software Schindelin et al. (2012) for two campaigns. We used the particle data from Fig. 3(b) for 145 successive frames and a particle tracking algorithm to gather particle velocities (in mm/s). Determining the accurate average interparticle separation, $(a)$, for dust clouds composed of particles of varying sizes and experiencing different spatial plasma environments is challenging. However, we calculated “$a$” from regions near vortex motion with visually no flow but only thermal motion. We obtained an estimated value of $a\approx 224~{}\mu$m and a typical dust density, $n_{d}\approx 2.3\times 10^{4}$ cm-3 in the diffused plasma region. Figure 3(b) depicts one such particle arrangement (from Campaign V) we used to estimate $a$.
@@
Dust particles typically have a charge between [2 - 5] ×$\times 10^{4}$$e$. The floating potential on a spherical dust particle is the basis for the collisionless orbital motion limited model, which is used to compute the charge. We fitted the theoretical Gaussian velocity distribution to the Probability Distribution Function (PDF) of the particle velocity data (averaged PDF from 145 frames), which we plotted in Fig. 3(a). A theoretical fitting in comparison with experimental data yields dust temperature ranges of 0.6 - 2.2 eV. We see that in a three-dimensional situation, the dust temperature will vary since it is computed using a two-dimensional velocity field. For two datasets (II and VI), we additionally computed the 2D radial distribution function (RDF). The dynamics are thermal, and there is no bulk flow in the area used for this computation. With a tiny initial peak, the RDF in Fig. 4 illustrates the equilibrium dust dynamics in a somewhat liquid regime.